Circular No. 6861

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Garnavich, Center for Astrophysics, reports that the High-Z Super-

nova Search Team (IAUC 6160, 6646) has discovered 11 supernovae on CCD images taken with the Cerro Tololo 4-m telescope by N. Suntzeff, M. Phillips, R. C. Smith, A. Clocchiatti, A. Diercks, P. Challis, and B. Schmidt. The supernovae were identified by subtracting images taken on Jan. 23–24 and Feb. 26–27 UT from those obtained Mar. 25–26, and confirmed with spectra and images taken on Mar. 5, 26, and 27 with the Keck-2 telescope by A. V. Filippenko, A. G. Riess, D. C. Leonard, and E. C. Moran.

SN  1998 UT  α_{2000}  δ_{2000}  I  z type
1998ac Feb. 27 9^{h}31^{m}20.^{s}00 − 5^{o}16′24′′.3 23.7 0.46 Ia
1998ad Mar. 25 9 29 55.71 − 5 21 59.3 23.4 0.44 II?
1998ae Mar. 25 9 30 47.53 − 4 38 51.0 23.9 1.1  ?
1998af Mar. 25 9 30 36.98 − 5 04 51.1 23.7 0.23 II
1998ag Mar. 25 13 43 36.42 + 0 08 02.2 23.5 0.67 Ia
1998ah Mar. 26 9 28 15.74 − 5 21 57.9 22.6 0.43 Ia
1998ai Mar. 26 13 42 27.59 − 0 06 47.2 22.7 0.49 Ia
1998aj Mar. 26 9 27 55.76 − 4 59 56.9 24.1 0.83 Ia
1998ak Mar. 26 13 44 16.22 − 0 10 49.3 23.7 0.74 Ia?
1998al Mar. 26 13 43 05.37 − 0 25 43.0 23.4 0.71 Ia
1998am Mar. 26 13 41 21.38 − 0 32 33.8 23.8 0.24 II

The supernovae are within 2′′ of their host galaxies. Finder charts may be requested by sending electronic mail to pgarnavich@cfa.harvard.edu.

1SAX J0544.1–710 G. Cusumano, T. Mineo, and L. Nicastro, Istituto di Fisica Cosmica

ed Applicazioni all’Informatica, CNR, Palermo; and G. L. Israel, Osservatorio Astronomico di Roma, report the discovery of a new x-ray pulsar, designated as 1SAX J0544.1–710, during a BeppoSAX observation of the Large Magellanic Cloud on 1996 Oct. 27–28: “The source position is α = 5^{h}44^{m}06^{s}, δ = −71^{o}00′.7 (equinox 2000.0; estimated error radius 40′′). The source shows pulsations at a period of 96.079 ± 0.007 s with a nearly sinusoidal pulse profile. The energy spectrum is well modelled by an absorbed power law with photon index 0.52 ± 0.08. The unabsorbed flux is 2.9 x 10^{−12} erg s^{−1} cm^{−2} in the energy band 2-10 keV. A public ASCA observation performed on 1996 Sept. 6 confirms the presence of the 96-s pulsations. ROSAT and Einstein data suggest the transient nature of the source.”

1998 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green